OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution Equation #1 7x - 14 = 6x + 2 Equation #2 7 - 2x = 5x + 21 A) The solution to equation #1 is smaller than the solution to equation #2. B) The solution to equation #1 is larger than the solution to equation #2. C) The solution to equation #1 is the same as the solution to equation #2. D) None of the statements above describe solutions to equations shown.

jimthompson5910 (jim_thompson5910):

Solving equation # 1 7x-14=6x+2 7x=6x+2+14 7x-6x=2+14 x=2+14 x=16 So the solution to equation #1 is x = 16 ------------------------------------------------------- Solving equation # 2 7-2x=5x+21 -2x=5x+21-7 -2x-5x=21-7 -7x=21-7 -7x=14 x=14/(-7) x = -2 So the solution to equation #2 is x = -2 ================================================ ================================================ In short, the solutions to equations #1 and #2 are x = 16 and x = -2 respectively. Therefore, the solution to equation #1 is larger than the solution to equation #2 So the answer is choice B

OpenStudy (anonymous):

You are a life savior, thank you!:))) Plus that made a lot of sense.

jimthompson5910 (jim_thompson5910):

You're welcome. Glad everything is clicking and that I could be of help.